Use the method of disks/rings to determine the volume of the solid obtained by rotating the region bounded by y=2x^2 and y=x^3 about the x-axis.


1
Expert's answer
2022-05-20T08:04:24-0400
"2x^2=x^3"

"x_1=0, x_2=2"

"V=\pi\displaystyle\int_{0}^{2}\big((2x^2)^2-(x^3)^2\big)dx"

"=\pi\big[\dfrac{4x^5}{5}-\dfrac{x^7}{7}\big]\begin{matrix}
 2 \\
 0
\end{matrix}=\pi(\dfrac{4(2)^5}{5}-\dfrac{(2)^7}{7}-0)"

"=\dfrac{256\pi}{35}({units}^3)"